OpenStudy (anonymous):

medals how many solutions does this system have 4x + 16y = 12, y = -1/4x + 9/12

OpenStudy (anonymous):

4x + 16y = 12 --- reduces to x + 4y = 3 y = -1/4x + 9/12 1/4x + y = 3/4 -- multiply by 4) x + 4y = 3 they are the same line...therefore, they have infinite solutions